OpenStudy (anonymous):

What is the sum of a 6–term geometric series if the first term is 22 and the last term is 369,754?

OpenStudy (anonymous):

22+22*7+22*7^2+22*7^3+22*7^4+22*7^5 = 431376

OpenStudy (anonymous):

very nice. where did the 7 come from?

OpenStudy (anonymous):

you know the 6th term is 369754. so 22x^5=369754 x^5 = 16807 x = 16807^(1/5) x = 7